# Visibility-Aware RRT* for Safety-Critical Navigation of Perception-Limited Robots in Unknown Environments

### v1

@@ Abstract
Safe autonomous navigation in unknown environments remains a critical challenge for robots with limited sensing capabilities. While safety-critical control techniques, such as Control Barrier Functions (CBFs), have been proposed to ensure safety, their effectiveness relies on the assumption that the robot has complete knowledge of its surroundings. In reality, robots often operate with restricted field-of-view and finite sensing range, which can lead to collisions with unknown obstacles if the planning algorithm is agnostic to these limitations. To address this issue, we introduce the visibility-aware RRT* algorithm that combines sampling-based planning with CBFs to generate safe and efficient global reference paths in partially unknown environments. The algorithm incorporates a collision avoidance CBF and a novel visibility CBF, which guarantees that the robot remains within locally collision-free regions, enabling timely detection and avoidance of unknown obstacles. We conduct extensive experiments interfacing the path planners with two different safety-critical controllers, wherein our method outperforms all other compared baselines across both safety and efficiency aspects.  [^1] Our project page can be found at: https://www.taekyung.me/visibility-rrt
@@ I INTRODUCTION
To ensure safe navigation, there has been growing interest in safety-critical controllers, such as those based on Control Barrier Functions (CBFs) [2], which enforce safety objectives on controlled systems. These controllers theoretically guarantee safety by keeping the robot within a safe set, assuming that nearby obstacles are fully known a priori. However, in practice, the effectiveness of these controllers relies heavily on the robot’s ability to perceive potential obstacles in the local environment. If the global planner does not account for the robot’s limited sensing capabilities under the presence of unknown obstacles, the theoretical safety guarantees may not hold, leading to potentially hazardous situations.
@@
Several approaches have been proposed to improve the safety of navigation under sensing limitations. Some methods use nonlinear optimization to generate [3] or refine [4] perception-aware paths, but these approaches can be computationally intensive and the solution convergence is not guaranteed. Other approaches first plan a path and then optimize the heading angle to enhance safety based on sensing objectives [5, 6, 7], although they do not always guarantee safety. Another method computes local safe regions during the exploration of candidate paths to account for local sensing information and ensure safety [8], but it requires storing all combinations of such regions, making the planning module overly computationally expensive.
@@
In this paper, we propose a visibility-aware RRT* algorithm that explicitly reasons about the robot’s limited sensing capabilities and combines the benefits of sampling-based planning with CBFs to enable safe and efficient navigation in partially unknown environments. Our approach incorporates two types of CBFs: (i) a collision avoidance CBF to ensure that the planned path is collision-free with respect to known obstacles, and (ii) a novel visibility CBF, which we introduce in this work, that guarantees the local tracking controller following the resulting path will always keep the robot within locally collision-free regions. This ensures that the robot can detect unknown obstacles in a timely manner, enabling the robot to avoid them by virtue of the local controller. We demonstrate the effectiveness of our method through extensive simulations and comparisons with other planning methods. The results indicate that our method can be seamlessly incorporated with safety-critical controllers, outperforming existing baselines in both safety and efficiency metrics across multiple scenarios.
@@ II-1 System Description
II-2 Linear Quadratic Regulator
Linear Quadratic Regulator (LQR) computes optimal control policies for linear systems, and can be applied to nonlinear systems by linearizing the system around an operating point $({\bm{x}}_{o},{\bm{u}}_{o})$, i.e., taking the first-order Taylor expansion of a nonlinear dynamics $F$ around this point:
$\dot{{\bm{x}}}\approx F({\bm{x}}_{o},{\bm{u}}_{o})+A\bar{{\bm{x}}}+B\bar{{\bm{u}}},$ (2)
where $\bar{{\bm{x}}}={\bm{x}}-{\bm{x}}_{o}$, $\bar{{\bm{u}}}={\bm{u}}-{\bm{u}}_{o}$, $A=\frac{\partial F}{\partial{\bm{x}}}|_{({\bm{x}}_{o},{\bm{u}}_{o})}$, $B=\frac{\partial F}{\partial{\bm{u}}}|_{({\bm{x}}_{o},{\bm{u}}_{o})}$. Given an infinite time horizon cost function:
$J=\int_{0}^{\infty}(\,\bar{{\bm{x}}}^{\top}Q\bar{{\bm{x}}}+\bar{{\bm{u}}}^{\top}R\bar{{\bm{u}}}\,)\,\mathrm{d}t,$ (3)
where $Q=Q^{\top}\succeq 0$ and $R=R^{\top}\succ 0$ are weight matrices, one can compute the locally optimal control policy $\mathcal{K}_{\text{lqr}}^{*}({\bm{x}})=-K_{\text{lqr}}{\bm{x}}$ by solving the algebraic Riccati equation:
$A^{\top}P+PA-PBR^{-1}B^{\top}P+Q=0.$ (4)
Solving (4) for $P$, the LQR gain matrix is:
$K_{\text{lqr}}=R^{-1}B^{\top}P.$ (5)
II-3 Control Barrier Functions
Definition 1 (Control Barrier Function (CBF)).
@@ II-3 Control Barrier Functions
for all ${\bm{x}}\in\mathcal{X}$ and $L_{f}h({\bm{x}})$ and $L_{g}h({\bm{x}})$ denote the Lie derivatives of $h$ along $f$ and $g$, respectively [9].
@@
Given a CBF $h$ with the associated set $\mathcal{C}$, any Lipschitz continuous controller ${\bm{u}}\in\mathcal{K}_{\textup{cbf}}({\bm{x}})$, with $\mathcal{K}_{\textup{cbf}}({\bm{x}})=\{{\bm{u}}\in\mathcal{U}:L_{f}h({\bm{x}})+L_{g}h({\bm{x}}){\bm{u}}+\alpha(h({\bm{x}}))\geq 0\}$, renders $\mathcal{C}$ forward invariant for System (1). See [9].
II-4 High Order CBF
The concept of CBFs has been extended to High Order CBFs (HOCBFs) [10], which provide a more general formulation that can be used for constraints of high relative degree. In this work, we focus on time-invariant HOCBFs, which generalize the concept of exponential CBFs proposed in [11].
For a $r_{b}^{\textup{th}}$ order differentiable function $h:\mathcal{X}\rightarrow\mathbb{R}$, we define a series of functions $\psi_{i}:\mathcal{X}\rightarrow\mathbb{R}$, $i=0,\ldots,r_{b}$, as:
$\psi_{0}({\bm{x}})$ $:=h({\bm{x}}),$ (7a)
$\psi_{1}({\bm{x}})$ $:=\dot{\psi}_{0}({\bm{x}})+\alpha_{1}(\psi_{0}({\bm{x}})),$ (7b)
@@ II-4 High Order CBF
$\psi_{r_{b}}({\bm{x}})$ $:=\dot{\psi}_{r_{b}-1}({\bm{x}})+\alpha_{r_{b}}(\psi_{r_{b}-1}({\bm{x}})),$ (7c)
where $\alpha_{i}:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$, $i=1,\ldots,r_{b}$, denote class $\mathcal{K}$ functions, and $\dot{\psi}_{i}({\bm{x}})=L_{f}\psi_{i}({\bm{x}})+L_{g}\psi_{i}({\bm{x}}){\bm{u}}$. We further define a series of sets $\mathcal{C}_{i}$ associated with these functions as:
$\mathcal{C}_{1}$ :=$:=\{{\bm{x}}\in\mathcal{X}:\psi_{0}({\bm{x}})\geq 0\},$ (8a)
$\mathcal{C}_{2}$ :=$:=\{{\bm{x}}\in\mathcal{X}:\psi_{1}({\bm{x}})\geq 0\},$ (8b)
@@
$\mathcal{C}_{r_{b}}$ :=$:=\{{\bm{x}}\in\mathcal{X}:\psi_{r_{b}-1}({\bm{x}})\geq 0\}.$ (8c)
Definition 2 (HOCBF).
The function $h:\mathcal{X}\rightarrow\mathbb{R}$ is a HOCBF of relative degree $r_{b}$ for System (1) if there exist differentiable class $\mathcal{K}$ functions $\alpha_{1},\alpha_{2},\ldots,\alpha_{r_{b}}$ such that
$\psi_{r_{b}}({\bm{x}})\geq 0,\quad\forall{\bm{x}}\in\mathcal{C}_{1}\cap\mathcal{C}_{2}\cap\ldots\cap\mathcal{C}_{r_{b}}.$ (9)
@@
Given a HOCBF $h$ with the associated sets $\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{r_{b}}$, if ${\bm{x}}(t_{0})\in\mathcal{C}_{1}\cap\mathcal{C}_{2}\cap\ldots\cap\mathcal{C}_{r_{b}}$, then any Lipschitz continuous controller ${\bm{u}}(t)\in\mathcal{K}_{\textup{hocbf}}({\bm{x}})$, with $\mathcal{K}_{\textup{hocbf}}({\bm{x}})=\{{\bm{u}}\in\mathcal{U}:\psi_{r_{b}}({\bm{x}})\geq 0\}$, renders the set $\mathcal{C}_{1}\cap\mathcal{C}_{2}\cap\ldots\cap\mathcal{C}_{r_{b}}$ forward invariant for system (1) [10].
@@ III PROBLEM FORMULATION
Consider a robot modeled as the nonlinear system (1) navigating in an unknown environment $\mathcal{W}$. The environment $\mathcal{W}$ consists of two types of obstacles: a set $\mathcal{O}_{0}\subset\mathcal{W}$ of known obstacles at the initial time $t_{0}$, and a set $\mathcal{H}\subset\mathcal{W}$ of unknown (or “hidden”) obstacles that can be sensed by the robot on-the-fly as it navigates through the environment. The true obstacle-free space $\mathcal{S}$ is then defined as $\mathcal{S}=\mathcal{W}\setminus(\mathcal{O}_{0}\cup\mathcal{H})$. A local tracking control policy $\pi:[t_{0},\infty)\times\mathcal{X}\times\mathcal{X}\to\mathcal{U}$ takes the current time $t$, the current state ${\bm{x}}_{t}:=\bm{x}(t)$, and a reference state ${\bm{x}}_{t}^{\textup{ref}}\in\mathcal{X}$ as inputs, and generates as outputs the control inputs that track a nominal path from the current to reference state. Given the initial condition ${\bm{x}}(t_{0})={\bm{x}}_{0}$, $t\geq t_{0}$, the robot dynamics obey:
$\dot{{\bm{x}}}(t)=f({\bm{x}}(t))+g({\bm{x}}(t))\pi(t,{\bm{x}}_{t},{\bm{x}}_{t}^{\textup{ref}}),\quad{\bm{x}}(t_{0})={\bm{x}}_{0}.$ (10)
The navigation problem is divided into two hierarchical steps: 1) finding a global reference path $p^{\textup{ref}}$, and 2) tracking $p^{\textup{ref}}$ using a local controller.
Global Path Planning: Given a goal state ${\bm{x}}_{\textup{goal}}\in\mathcal{X}$, the objective of the global path planner is to find a global reference path $p^{\textup{ref}}:\mathcal{X}\to\mathcal{P}$, represented by a set of waypoints $\mathcal{P}=\{{\bm{x}}_{i}^{\textup{ref}}\in\mathcal{X}\}_{i=0}^{N}$, where ${\bm{x}}_{0}^{\textup{ref}}={\bm{x}}_{0}$ and ${\bm{x}}_{N}^{\textup{ref}}={\bm{x}}_{\textup{goal}}$ are the initial and final waypoints in $\mathcal{P}$, respectively. For any given state ${\bm{x}}_{t}$, the function yields the next waypoint that the robot should navigate towards as:
${\bm{x}}_{i}^{\textup{ref}}=p^{\textup{ref}}({\bm{x}}_{t}),$ (11)
where ${\bm{x}}_{i}^{\textup{ref}}\in\mathcal{P}\subset\mathcal{X}$ indicates the immediate waypoint for navigation. The reference path should satisfy:
@@
We define the known collision-free set at $t_{0}$ as $\mathcal{S}_{0}=\mathcal{W}\setminus\mathcal{O}_{0}$. A reference path $p^{\textup{ref}}$ is considered collision-free if it satisfies the condition:
@@
$\mathcal{F}_{t}=\bigg{\{}{\bm{p}}\in\mathcal{S}\mid$ $\,\|{\bm{p}}-{\bm{r}}_{t}\|\leq l_{\textup{range}}$
$\cap\,|\,\angle({\bm{p}}-{\bm{r}}_{t},\theta_{t})\,|\leq\frac{\theta_{\textup{fov}}}{2}\bigg{\}},$ (14)
where $\|{\bm{p}}-{\bm{r}}_{t}\|$ represents the Euclidean distance between a point ${\bm{p}}\in\mathcal{W}$ and the robot’s position ${\bm{r}}_{t}$, $\angle({\bm{p}}-{\bm{r}}_{t},\theta_{t})$ denotes the angle between the vector ${\bm{p}}-{\bm{r}}_{t}$ and the robot’s orientation $\theta_{t}$. Given the accumulated sensory information collected from time $t_{0}$ to $t$, we define the known collision-free space $\mathcal{B}_{t}$ at time $t$ as:
$\mathcal{B}_{t}:=\bigcup_{\tau=t_{0}}^{t}\mathcal{F}_{\tau}\subseteq\mathcal{S}.$ (15)
Consequently, the part of the environment that has not been sensed yet is $\mathcal{S}\setminus\mathcal{B}_{t}$, and is called the unmeasured space at time $t$.
It is important to note that the collision-free path criterion (13) treats all unmeasured space $\mathcal{S}\setminus\mathcal{B}_{t_{0}}$ at time $t_{0}$ as free space. However, while following the path $p^{\textup{ref}}$ using the local controller $\pi$ after the planning cycle, unknown obstacles $\mathcal{H}$ might exist within such unmeasured spaces.
@@
A reference path $p^{\textup{ref}}$ is called visibility-aware w.r.t. the dynamics (1) if the updated state ${\bm{x}}(t+\Delta t)$, for a small time increment $\Delta t>0$, under the local tracking controller $\pi(t,{\bm{x}}_{t},p^{\textup{ref}}({\bm{x}}_{t}))$ satisfies:
${\bm{x}}(t+\Delta t)\in\mathcal{B}_{t},\quad\forall t\geq t_{0},\quad\forall{\bm{x}}_{t}\in\mathcal{X}.$ (16)
In the following sections, we present the construction of a global path planner that imposes (16) (see Sec. IV), and demonstrate how (16) ensures hidden obstacles are avoidable, in the sense that the local tracking controller is able to generate collision-free trajectories (see Sec. V).
Local Tracking Control: A local control policy $\pi$ is used to stabilize the systems around the subsequent waypoint ${\bm{x}}_{i}^{\textup{ref}}$ (11) and to ensure that the robot remains within the true collision-free set $\mathcal{S}$. The local controller is assumed to be capable of tracking the path between subsequent waypoints with a tracking error bounded by a maximum displacement $\epsilon$. As $p^{\textup{ref}}$ continuously updates the waypoint, it drives the robot towards the goal state ${\bm{x}}_{\textup{goal}}$. There exists a plethora of methodologies to design such controllers [9, 12]. Since the primary focus of this paper is on designing a visibility-aware path planner, we will discuss two case studies of such controllers in Sec. V and Sec. VI to highlight the advantages of our proposed planner when utilized alongside various local controllers.

### v2

@@ Abstract
Safe autonomous navigation in unknown environments remains a critical challenge for robots with limited sensing capabilities. While safety-critical control techniques, such as Control Barrier Functions (CBFs), have been proposed to ensure safety, their effectiveness relies on the assumption that the robot has complete knowledge of its surroundings. In reality, robots often operate with restricted field-of-view and finite sensing range, which can lead to collisions with unknown obstacles if the planner is agnostic to these limitations. To address this issue, we introduce the Visibility-Aware RRT* algorithm that combines sampling-based planning with CBFs to generate safe and efficient global reference paths in partially unknown environments. The algorithm incorporates a collision avoidance CBF and a novel visibility CBF, which guarantees that the robot remains within locally collision-free regions, enabling timely detection and avoidance of unknown obstacles. We conduct extensive experiments interfacing the path planners with two different safety-critical controllers, wherein our method outperforms all other compared baselines across both safety and efficiency aspects. [Project Page] [^1] Project page: https://www.taekyung.me/visibility-rrt [Code] [Video]
@@ I INTRODUCTION
Figure 1: In this paper we develop an RRT*-based planner that accounts for sensing limitations, and in conjunction with a tracking controller, guarantees avoidance of a priori unknown obstacles. This figure shows trajectories generated by baseline (Top) and visibility-aware (Bottom, Ours) planning methods, tracked by the rover using a CBF-QP controller (baselines and results presented in Section VI.). Inset: Snapshot of the Signed Distance Field map just before the collision while following the baseline trajectory, showing that the unknown obstacle (which is located in the dashed circle, but not in the FOV of the robot) is not detected in time, leading to collision.
@@
To ensure safe navigation, there has been growing interest in safety-critical controllers, such as those based on Control Barrier Functions (CBFs) [2], which enforce safety objectives on controlled systems. These controllers theoretically guarantee safety by keeping the robot within a safe set, assuming that nearby obstacles are fully known a priori. However, in practice, the effectiveness of these controllers relies heavily on the robot’s ability to perceive obstacles in the local environment. If the global planner does not account for the robot’s limited sensing capabilities under the presence of unknown obstacles, the theoretical safety guarantees may not hold, leading to possible safety violations.
@@
Several approaches have been proposed to improve the safety of navigation under sensing limitations. Some methods use nonlinear optimization to generate [3] or refine [4] perception-aware paths, but these approaches can be computationally intensive and the solution convergence is not guaranteed. Other approaches first plan a path and then optimize the heading angle to enhance safety based on sensing objectives [5, 6, 7], although they do not always guarantee safety. Another method computes local safe regions during the exploration of candidate paths to account for local sensing information and ensure safety [8], but it requires storing all combinations of such regions, making the planning module overly computationally expensive. Additionally, belief-space planning approaches [9, 10] explicitly model state-estimation uncertainty within sampling-based frameworks. However, these approaches primarily focus on sensing uncertainty rather than limited field-of-view, which is the focus of our work.
@@
In this paper, we propose the Visibility-Aware RRT* algorithm that explicitly reasons about the robot’s limited sensing capabilities and combines the benefits of sampling-based planning with CBFs to enable safe and efficient navigation in partially unknown environments. Our approach incorporates two types of CBFs: (i) a collision avoidance CBF to ensure that the planned path is collision-free with respect to known obstacles, and (ii) a novel visibility CBF, which we introduce in this work, that guarantees the local tracking controller following the resulting path will always keep the robot within locally collision-free regions. This ensures that the robot can detect unknown obstacles in a timely manner, enabling the robot to avoid them by virtue of the local controller. We demonstrate the effectiveness of our method through extensive simulations and experiments (see Fig. 1), comparing it with other planning methods. The results demonstrate that our approach effectively addresses the challenges faced by both CBF-based controller [2] and Gatekeeper algorithm [11], outperforming existing baselines across multiple scenarios.
@@ II-1 System Description
II-2 Control Barrier Functions
Definition 1 (Control Barrier Function (CBF) [2]).
@@ II-2 Control Barrier Functions
for all ${\bm{x}}\in\mathcal{X}$ and $L_{f}h({\bm{x}})$ and $L_{g}h({\bm{x}})$ denote the Lie derivatives of $h$ along $f$ and $g$, respectively.
@@
[2] Given a CBF $h$ with the associated set $\mathcal{C}$, any Lipschitz continuous controller ${\bm{u}}\in\mathcal{K}_{\textup{cbf}}({\bm{x}})$, with $\mathcal{K}_{\textup{cbf}}({\bm{x}})\coloneqq\{{\bm{u}}\in\mathcal{U}:L_{f}h({\bm{x}})+L_{g}h({\bm{x}}){\bm{u}}+\alpha(h({\bm{x}}))\geq 0\}$, renders $\mathcal{C}$ forward invariant for System (1).
II-3 High Order CBF
The concept of CBFs has been extended to High Order CBFs (HOCBFs) [12], which provide a more general formulation that can be used for constraints of high relative degree.
For a $r^{\textup{th}}$ order differentiable function $h:\mathcal{X}\rightarrow\mathbb{R}$, we define a series of functions $\psi_{i}:\mathcal{X}\rightarrow\mathbb{R}$, $i=0,\ldots,r$, as:
$\psi_{0}({\bm{x}})$ $\coloneqq h({\bm{x}}),$ (3a)
$\psi_{1}({\bm{x}})$ $\coloneqq\dot{\psi}_{0}({\bm{x}})+\alpha_{1}(\psi_{0}({\bm{x}})),$ (3b)
@@ II-3 High Order CBF
$\psi_{r}({\bm{x}})$ $\coloneqq\dot{\psi}_{r-1}({\bm{x}})+\alpha_{r}(\psi_{r-1}({\bm{x}})),$ (3c)
where $\alpha_{i}:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$, $i=1,\ldots,r$, denote class $\mathcal{K}$ functions, and $\dot{\psi}_{i}({\bm{x}})=L_{f}\psi_{i}({\bm{x}})+L_{g}\psi_{i}({\bm{x}}){\bm{u}}$. We further define a series of sets $\mathcal{C}_{i}$ associated with these functions as:
$\mathcal{C}_{1}$$\coloneqq\{{\bm{x}}\in\mathcal{X}:\psi_{0}({\bm{x}})\geq 0\},$ (4a)
$\mathcal{C}_{2}$$\coloneqq\{{\bm{x}}\in\mathcal{X}:\psi_{1}({\bm{x}})\geq 0\},$ (4b)
@@
$\mathcal{C}_{r}$$\coloneqq\{{\bm{x}}\in\mathcal{X}:\psi_{r-1}({\bm{x}})\geq 0\}.$ (4c)
Definition 2 (HOCBF [12]).
The function $h:\mathcal{X}\rightarrow\mathbb{R}$ is a HOCBF of relative degree $r$ for System (1) if there exist differentiable class $\mathcal{K}$ functions $\alpha_{1},\alpha_{2},\ldots,\alpha_{r}$ such that
$\psi_{r}({\bm{x}})\geq 0,\quad\forall{\bm{x}}\in\mathcal{C}_{1}\cap\mathcal{C}_{2}\cap\ldots\cap\mathcal{C}_{r}.$ (5)
@@
[12] Given a HOCBF $h$ with the associated sets $\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{r}$, if ${\bm{x}}(t_{0})\in\mathcal{C}_{1}\cap\mathcal{C}_{2}\cap\ldots\cap\mathcal{C}_{r}$, then any Lipschitz continuous controller ${\bm{u}}(t)\in\mathcal{K}_{\textup{hocbf}}({\bm{x}})$, with $\mathcal{K}_{\textup{hocbf}}({\bm{x}})\coloneqq\{{\bm{u}}\in\mathcal{U}:\psi_{r}({\bm{x}})\geq 0\}$, renders the set $\mathcal{C}_{1}\cap\mathcal{C}_{2}\cap\ldots\cap\mathcal{C}_{r}$ forward invariant for System (1).
@@ III PROBLEM FORMULATION
Consider a robot modeled as the nonlinear system (1) navigating in an unknown environment $\mathcal{W}$. The environment $\mathcal{W}$ consists of two types of obstacles: a set $\mathcal{O}_{0}\subset\mathcal{W}$ of known obstacles at the initial time $t_{0}$, and a set $\mathcal{H}\subset\mathcal{W}$ of unknown (or “hidden”) obstacles that can be sensed by the robot on-the-fly as it navigates through the environment. The true obstacle-free space $\mathcal{S}$ is then defined as $\mathcal{S}=\mathcal{W}\setminus(\mathcal{O}_{0}\cup\mathcal{H})$. A local tracking control policy $\pi$ generates control inputs that track a nominal path from the current to reference state.
The navigation problem is divided into two hierarchical steps: 1) finding a global reference path $p^{\textup{ref}}$, and 2) tracking $p^{\textup{ref}}$ using a local controller $\pi$.
Global Path Planning: Given a goal state ${\bm{x}}_{\textup{goal}}\in\mathcal{X}$, the objective of the global path planner is to find a global reference path $p^{\textup{ref}}:\mathcal{X}\to\mathcal{P}$, represented by a set of waypoints $\mathcal{P}=\{{\bm{x}}_{j}^{\textup{ref}}\in\mathcal{X}\}_{j=0}^{N}$, where ${\bm{x}}_{0}^{\textup{ref}}={\bm{x}}_{0}$ and ${\bm{x}}_{N}^{\textup{ref}}={\bm{x}}_{\textup{goal}}$ are the initial and final waypoints in $\mathcal{P}$, respectively. For any given state ${\bm{x}}_{t}$, the function yields the next waypoint that the robot should navigate towards as:
${\bm{x}}_{j}^{\textup{ref}}=p^{\textup{ref}}({\bm{x}}_{t}),$ (6)
where the index $j$ increases monotonically with time, and ${\bm{x}}_{j}^{\textup{ref}}\in\mathcal{P}\subset\mathcal{X}$ indicates the immediate waypoint for navigation. The reference path should satisfy:
@@
We define the known collision-free set at $t_{0}$ as $\mathcal{S}_{0}\coloneqq\mathcal{W}\setminus\mathcal{O}_{0}$. A reference path $p^{\textup{ref}}$ is considered collision-free if it satisfies the condition:
@@
[…]
$|\,\angle({\bm{p}}-{\bm{r}}_{t},\theta_{t})\,|\leq\frac{\theta_{\textup{fov}}}{2}\,\cap\,\textup{LoS}({\bm{r}}_{t},{\bm{p}})\bigg{\}},$ (9)
where $\angle({\bm{p}}-{\bm{r}}_{t},\theta_{t})$ denotes the angle between the vector ${\bm{p}}-{\bm{r}}_{t}$ and the robot’s orientation $\theta_{t}$, and $\textup{LoS}({\bm{r}}_{t},{\bm{p}})$ indicates that the line of sight from ${\bm{r}}_{t}$ to ${\bm{p}}$ does not intersect any obstacles; i.e., $\forall s\in[0,1],\ {\bm{r}}_{t}+s({\bm{p}}-{\bm{r}}_{t})\in\mathcal{S}$.
Given the accumulated sensory information collected from $t_{0}$ to $t$, we define the known collision-free space $\mathcal{B}_{t}$ at $t$ as:
$\mathcal{B}_{t}\coloneqq\bigcup_{\tau=t_{0}}^{t}\mathcal{F}_{\tau}\subseteq\mathcal{S}.$ (10)
Consequently, the part of the environment that has not been sensed yet is $\mathcal{S}\setminus\mathcal{B}_{t}$, called the unmeasured space at $t$.
It is important to note that the collision-free path criterion (8) treats all unmeasured space $\mathcal{S}\setminus\mathcal{B}_{t_{0}}$ at $t_{0}$ as free space. However, while following the path $p^{\textup{ref}}$ using the local controller $\pi$ after the planning cycle, unknown obstacles $\mathcal{H}$ might exist within such unmeasured spaces.
@@
A reference path $p^{\textup{ref}}$ is called visibility-aware w.r.t. the dynamics (1) if under the local tracking controller $\pi$, the updated state ${\bm{x}}(t+\Delta t)$ satisfies:
${\bm{x}}(t+\Delta t)\in\mathcal{B}_{t},\quad\forall t\geq t_{0},\quad\forall{\bm{x}}_{t}\in\mathcal{X},$ (11)
where $\Delta t>0$ is a small time increment.
We formally define the problem that our global path planner aims to solve:
Problem 1.
Given a path planning problem ($\mathcal{S}_{0},{\bm{x}}_{0},{\bm{x}}_{\textup{goal}}$) for System (1), find a path $p^{\textup{ref}}$ such that for all $j=0,\ldots,N-1$, the state and input trajectories denoted as $\bm{\sigma}_{j}:=({\bm{x}}(\tau),{\bm{u}}(\tau))$ connecting consecutive waypoints ${\bm{x}}^{\textup{ref}}_{j}$ and ${\bm{x}}^{\textup{ref}}_{j+1}$ during the time interval $[t_{j},t_{j+1})$ satisfies ${\bm{x}}(\tau)\in\mathcal{C}_{\textup{col}}\cap\mathcal{C}_{\textup{vis}}$ and ${\bm{u}}(\tau)\in\mathcal{K}_{\textup{col}}({\bm{x}}(\tau))\cap\mathcal{K}_{\textup{vis}}({\bm{x}}(\tau))$ for all $\tau\in[t_{j},t_{j+1})$. If such a path exists, return it; otherwise, report failure.
In the above formulation, $\mathcal{C}_{\textup{col}}$ and $\mathcal{C}_{\textup{vis}}$ are the sets that can be rendered forward invariant for System (1) under the two CBFs, which encode collision-avoidance (8) and visibility-awareness (11), respectively. The sets $\mathcal{K}_{\textup{col}}({\bm{x}})$ and $\mathcal{K}_{\textup{vis}}({\bm{x}})$ represent the admissible control inputs for System (1) that ensure these CBF constraints hold at each state ${\bm{x}}$. In the following sections, we detail how the global path planner imposes these two constraints (8), (11) in its search (see Sec. IV), and illustrate how these constraints ensure that any newly detected, previously unknown obstacles remain avoidable by the local tracking controller (see Sec. V).
Local Tracking Control: A local control policy $\pi$ is used to stabilize the systems around the subsequent waypoint ${\bm{x}}_{j}^{\textup{ref}}$ (6) and ensure that the robot remains within the true collision-free set $\mathcal{S}$. The local controller is assumed to be capable of tracking the path between subsequent waypoints with a tracking error bounded by a maximum displacement $\epsilon$. As $p^{\textup{ref}}$ continuously updates the waypoint, it drives the robot towards the goal state ${\bm{x}}_{\textup{goal}}$. There exist many methodologies to design such controllers [2, 11]. Since the primary focus of this paper is on designing a visibility-aware path planner, we will discuss two case studies of such controllers in Sec. V and Sec. VI to highlight the advantages of our proposed planner when utilized alongside various local controllers.